OpenStudy (anonymous):

I need a little help with my financial planning? Corrina purchased a motorcycle for $8,487. She made a down payment of $2,225. She applied for a four-year installment loan with an interest rate of 11.2%. What is the total cost of the motorcycle after four years? $7,797.60 $10,022.60 $10,568.64 $12,793.64

OpenStudy (anonymous):

8487 - 2225 = the financied amount

OpenStudy (anonymous):

then ia it simple interest or compound

OpenStudy (anonymous):

I'm pretty sure it's simple interest

OpenStudy (anonymous):

okay the its (8487 - 2225) get your answer do this answer times years times 0.112 = Interest paid 8487 plus interest paid. If its simple interest

OpenStudy (anonymous):

I got 8662.336 when i did that? I did 6262*(1/4)*0.112= 175.336+8487